OpenStudy (anonymous):

a carpet has perimeter 220 inches.the length of carpet is 22 inches mor than width.what is deimensions.Set problem to =0

OpenStudy (anonymous):

66 inches, 44 inches if the length is x and width is y, then perimeter= 2(x+y)=220 or, x+y=110 but, x=y+22 so, x+y=y+22+y=2y+22=110 from here, we can get, y=44 x=y+22=44+22=66